8 8
______ = ____
x + 6 9

solve the following equation. be sure to check each answer in the original equation if you multiply both side with an expression that contains the variable

Answer by checkley71(8403)

8/(X+6)=8/9 NOW CROSS MULTIPLY.
8(X+6)=8*9
8X+48=72
8X=72-48
8X=24
X=24/8
X=3 ANSWER.
PROOF
8/(3+6)=8/9
8/9=8/9